# Constructing a 3D Town from a Single Image

### v1

@@ Abstract
Acquiring detailed 3D scenes typically demands costly equipment, multi-view data, or labor-intensive modeling. Therefore, a lightweight alternative, generating complex 3D scenes from a single top-down image, plays an essential role in real-world applications. While recent 3D generative models have achieved remarkable results at the object level, their extension to full-scene generation often leads to inconsistent geometry, layout hallucinations, and low-quality meshes. In this work, we introduce 3DTown, a training-free framework designed to synthesize realistic and coherent 3D scenes from a single top-down view. Our method is grounded in two principles: region-based generation to improve image-to-3D alignment and resolution, and spatial-aware 3D inpainting to ensure global scene coherence and high-quality geometry generation. Specifically, we decompose the input image into overlapping regions and generate each using a pretrained 3D object generator, followed by a masked rectified flow inpainting process that fills in missing geometry while maintaining structural continuity. This modular design allows us to overcome resolution bottlenecks and preserve spatial structure without requiring 3D supervision or fine-tuning. Extensive experiments across diverse scenes show that 3DTown outperforms state-of-the-art baselines, including Trellis, Hunyuan3D-2, and TripoSG, in terms of geometry quality, spatial coherence, and texture fidelity. Our results demonstrate that high-quality 3D town generation is achievable from a single image using a principled, training-free approach. Project website: https://eric-ai-lab.github.io/3dtown.github.io/
Figure 1: 3D Scene Generation from a Single Image. Given a top-down reference image (center), 3DTown generates coherent and realistic 3D scenes that preserves geometry, texture, and layout compared to other state-of-the-art image-to-3D generation models. Our method also generalizes across diverse styles (right), producing high-quality outputs without any 3D training.
@@ 1 Introduction
Realistic 3D towns and structured environments serve as essential assets in simulation, robotics, digital content creation, and virtual world building. They enable scalable training for autonomous agents, immersive game environments, and rapid digital twin construction. However, constructing detailed and coherent 3D scenes typically requires either expensive 3D scanning equipment, multi-view data collection, or labor-intensive modeling. In contrast, generating 3D scenes from a single top-down image offers a lightweight and accessible alternative, making it possible to bootstrap realistic environments from minimal input.
@@
Despite its appeal, generating a coherent and realistic 3D scene from a single image presents several fundamental challenges. First, the synthesized scene must exhibit consistent geometry across novel views, which is difficult for pure volumetric rendering methods such as Neural Radiance Fields (NeRF) [35] or 3D Gaussian Splatting (3DGS) [18]. While these techniques excel at photorealistic appearance modeling, they often suffer from geometry artifacts, especially in occluded or sparsely visible regions, leading to multiview inconsistencies and structural implausibility [25, 5, 66, 64]. Second, the global layout of the generated scene must remain faithful to the input image. This is particularly difficult when considering the whole scene as one asset and using image-to-3D asset generators [59, 52, 24], which fail to preserve spatial relationships between elements, resulting in distorted or semantically misaligned arrangements. Third, the local fidelity of individual objects should align closely with the visual evidence in the input. Due to the resolution constraints of 3D representations and the domain shift from object-level training to scene-level inference, previous image-to-3D generators are prone to creating low-quality meshes and misaligning the texture.
@@
To address these challenges, we propose 3DTown, a training-free framework for generating complex 3D scenes from a single top-down image, by enhancing the capability of image-to-3D object generators. Our method combines two core components: region-based generation and spatial-aware 3D inpainting. Each targets specific challenges in existing pipelines. We divide the scene into overlapping regions and synthesize each independently. This modular approach enables spatial upscaling and improves local alignment by grounding generation on localized image crops. To maintain global coherence and object continuity across regions, we estimate a coarse 3D structure from monocular depth and landmark detection, forming a spatial prior. A masked rectified flow mechanism then completes missing parts while preserving known content, enhancing structural consistency and object-level fidelity throughout the scene.
@@
Through extensive experiments, we demonstrate that 3DTown generates realistic, diverse, and geometrically consistent 3D scenes from a single top-down image. Our method significantly outperforms strong baselines, including Trellis [59], Hunyuan3D-2 [52], and TripoSG [24], across both human preference and GPT-based evaluations. Quantitative results show notable gains in geometry quality, layout coherence, and texture fidelity, while qualitative comparisons highlight 3DTown’s ability to preserve spatial structure and fine-grained detail. These results underscore the effectiveness of our modular, training-free approach to 3D scene synthesis.
@@
  • We propose 3DTown, a training-free framework for generating structured 3D scenes from a single top-down image, leveraging pretrained object-centric generators for zero-shot scene asset synthesis.
@@
  • We conduct comprehensive evaluations on diverse scenes and show that 3DTown outperforms state-of-the-art baselines (Trellis, Hunyuan3D-2, and TripoSG) in geometry quality, layout coherence, and texture realism under both human and GPT-4o-based assessments.
@@ 2 Related Work
Recent progress in 2D generative modeling [42, 13, 4, 39] has inspired a growing body of work that combines image synthesis with external 3D reconstruction techniques to generate 3D scenes. A common strategy is to use pretrained image generation models to outpaint a given input view according to known or sampled camera trajectories, followed by reconstruction via depth fusion or volumetric representations such as Neural Radiance Fields (NeRF) [35] and 3D Gaussian Splatting (3DGS) [18]. Early approaches primarily targeted indoor scenes [56, 20, 14, 19], while more recent efforts have extended to arbitrary natural scenes [25, 1, 10, 5], achieving improved reconstruction quality through advanced depth fusion pipelines [64, 68, 66, 62, 45, 8, 63]. Despite producing visually compelling results, these methods often suffer from geometric inconsistencies caused by hallucinations during image generation, particularly in occluded or ambiguous regions.
To mitigate this, recent approaches explore panoramic [46, 23, 57, 44, 26] and multiview generation [29, 51, 11] to improve spatial coverage and structure. However, these approaches typically restrict the freedom of movement or require the camera motion for close views, and still struggle with accurate 3D geometry synthesis. The concurrent work Syncity [7] proposes block-wise generation via a 3D generator and assembles them to scenes, but all generated scenes have compact layouts and cannot handle the entire scene images. In contrast, we aim to generate geometry-consistent 3D assets with arbitrary layout directly from a single top-down image.
@@
Beyond 2D-to-3D approaches, recent research has explored the direct generation of 3D scenes using native 3D representations. One prominent line of work leverages large language models [9, 69, 2, 15, 47] or trained diffusion models [48, 28, 65, 54, 33] to synthesize scene layouts, which are then populated by retrieving or generating 3D assets according to the predicted configuration. While these methods enable semantically plausible arrangements, they are often constrained to predefined object categories and struggle to produce coherent background geometry. Furthermore, they frequently encounter inter-object collisions.
Meanwhile, a separate class of methods directly models the 3D scene generation process in latent 3D spaces [58, 34, 41, 31, 21, 3]. For instance, BlockFusion [58] trains a diffusion model to generate triplane representations for regional scene blocks, while LT3SD [34] and XCube [41] adopt truncated unsigned distance fields (TUDFs) or voxels to represent full scenes. Despite their architectural innovations, these models require large-scale, domain-specific datasets for training, such as indoor room datasets or urban layouts, which limits their generalizability to unseen scene types or out-of-distribution inputs. In contrast, our method avoids any training or fine-tuning and instead employs a modular, training-free pipeline that generates diverse 3D scenes from top-down images.
@@
Another line of research focuses on generating individual 3D assets from single images, aiming to reconstruct both geometry and appearance of the depicted objects. Early works [38, 27, 50, 30, 49] leverage 2D generative models in combination with NeRF or 3D Gaussian Splatting optimization to lift 2D images into 3D. However, these approaches often suffer from low geometry fidelity and slow optimization. To address these limitations, more recent methods [12, 61, 22, 53, 67] propose to directly generate 3D latent representations via diffusion models, while others [59, 52, 24] adopt rectified flow to improve both quality and efficiency. These methods typically rely on large-scale 3D datasets such as Objaverse-XL [6], which mainly consist of object-centric assets. As a result, applying them directly to full-scene synthesis introduces several challenges: limited 3D resolution becomes a bottleneck for encoding complex environments, and domain shift from object-level training to scene-level inference often leads to spatial inconsistencies and layout hallucinations.
In this work, we build upon pretrained 3D generation models [59] and address these limitations through a region-based generation strategy combined with spatial-aware 3D inpainting. This design enables us to produce 3D scene assets with both high geometric fidelity and global layout coherence from a single top-down image.
@@ 3 Method
Figure 2: Overview of the 3DTown Pipeline. Given a single top-down image, we first estimate a coarse scene structure via monocular depth and landmark extraction to initialize the scene latent (Spatial Prior Initialization). The scene is divided into overlapping regions for localized synthesis and progressively fused into a coherent global latent (Region-based Generation & Fusion). Each region is completed using a two-stage masked rectified flow pipeline with a sparse structure generator $\mathcal{G}_{s}$ and a structured latent generator $\mathcal{G}_{L}$ (Spatial-aware 3D Completion). The final 3D scene is decoded from the completed structured latent.
@@
From the scene image, we first construct the scene latents with structured latent representations (Sec. 3.1) from spatial priors (Sec. 3.2). Then, we divide the scene-level latent into region-level latents for sequential processing. For each region, we extract the latents from the latest scene-level structured latents and take those as priors for spatial-aware 3D completion (Sec. 3.3) by using the base 3D generator, Trellis [59]. Later, we fuse the updated region latents to the scene latents for cross-region consistency (Sec. 3.4). After finishing all regions, we leverage pretrained object decoders on the complete scene structured latents to obtain a 3D scene asset. Details are explained below.
@@ 3.1 Structured Latent Representation
To leverage the pretrained knowledge of the base 3D generator, we construct the scene with a structured latent representation $z$ [59]:
@@
For the general image-to-3D asset generation, there are two pretrained rectified flow transformers: a sparse structure generator $\mathcal{G}_{s}$ and a structured latent generator $\mathcal{G}_{L}$. At inference time, $\mathcal{G}_{s}$ first generates the active voxel positions $\{p_{i}\}^{L}$ from the noisy grids $V_{T}$ with Gaussian noise. These positions are then used by $\mathcal{G}_{L}$ to generate the corresponding latent features $\{f_{i}\}^{L}$ from noisy features $F_{T}$. Both generators are conditioned on an input image condition $C_{I}$, encoded by DINOv2 [37]. Finally, the structured latent representation $z$ is decoded into a 3D object $O$ using object decoders, which include different sparse 3D VAE decoders for 3D Gaussians, Radiance Fields, and mesh generation. The overall process is described as follows:
@@ 3.2 Spatial Prior Initialization
Given a top-down image, the image-to-3D generator may produce outputs in arbitrary orientations. To resolve this ambiguity and provide a consistent structural prior, we initialize the scene using point clouds. Specifically, we employ a monocular depth estimator [55] to predict a depth image and infer camera parameters, from which we construct pixel-wise point clouds. Due to occlusions, these point clouds contain missing regions, which will later be filled by the image-to-3D generator.
However, occluded areas of complex objects (e.g., buildings) may have multiple plausible completions. To enforce cross-region consistency in such cases, we propose first generating landmark objects independently and conditioning subsequent generations on their geometry. To achieve this, we use Florence2 [60] to propose landmark bounding boxes and SAM2 [40] to extract instance masks. Each detected landmark is then processed individually using the 3D generator to obtain instance-level meshes. These meshes are aligned with the raw point clouds using ICP [43], replacing the original landmark regions with mesh-derived point clouds.
@@ 3.3 Region-based Generation
While region-based generation improves local fidelity, it introduces a new challenge: how to maintain global consistency, especially across overlapping regions. Furthermore, since the 3D generator may create assets from any orientation, we need to alleviate the misalignment between image conditions and region latents to enable further region fusion. To address these challenges, we draw inspiration from training-free inpainting methods in 2D diffusion models, such as RePaint [32], and adapt a similar approach for 3D generation. We propose to use a masked rectified flow pipeline for 3D completion, which treats the partially completed global scene latent as a constraint and performs conditional generation over the current region by completing only the unknown parts.
@@
We initialize the latent variable $x_{T}\sim\mathcal{N}(0,I)$ with Gaussian noise, representing the unknown region at the final time step. For each timestep $t=T,T{-}1,\dots,1$, we perform $U$ resampling steps to improve stability and smoothness [32]. At each resampling iteration $u$, we first compute the flow field $v_{\theta}(x_{t},t)$ using the rectified flow model and apply an Euler update to obtain the intermediate latent:
@@
[…]

### v2

# Constructing A 3D Scene from a Single Image

@@ Abstract
Acquiring detailed 3D scenes typically demands costly equipment, multi-view data, or labor-intensive modeling. Therefore, a lightweight alternative, generating complex 3D scenes from a single top-down image, plays an essential role in real-world applications. While recent 3D generative models have achieved remarkable results at the object level, their extension to full-scene generation often leads to inconsistent geometry, layout hallucinations, and low-quality meshes. In this work, we introduce SceneFuse-3D, a training-free framework designed to synthesize coherent 3D scenes from a single top-down view. Our method is grounded in two principles: region-based generation to improve image-to-3D alignment and resolution, and spatial-aware 3D inpainting to ensure global scene coherence and high-quality geometry generation. Specifically, we decompose the input image into overlapping regions and generate each using a pretrained 3D object generator, followed by a masked rectified flow inpainting process that fills in missing geometry while maintaining structural continuity. This modular design allows us to overcome resolution bottlenecks and preserve spatial structure without requiring 3D supervision or fine-tuning. Extensive experiments across diverse scenes show that SceneFuse-3D outperforms state-of-the-art baselines, including Trellis, Hunyuan3D-2, TripoSG, and LGM, in terms of geometry quality, spatial coherence, and texture fidelity. Our results demonstrate that high-quality coherent 3D scene-level asset generation is achievable from a single top-down image using a principled, training-free pipeline. Project website: https://eric-ai-lab.github.io/3dtown.github.io/
Figure 1: 3D Scene Generation from a Single Image. Given a top-down reference image (center), SceneFuse-3D generates coherent and realistic 3D scenes that preserve geometry, texture, and layout compared to other state-of-the-art end-to-end image-to-3D generation models. Our method also generalizes across diverse styles (right), producing high-quality outputs without any 3D training.
@@ 1 Introduction
Constructing 3D scene environments serves as an essential component in simulation, robotics, digital content creation, and virtual world building. They enable scalable training for autonomous agents, immersive game environments, and rapid digital twin construction. However, constructing detailed and coherent 3D scenes typically requires either expensive 3D scanning equipment, multi-view data collection, or labor-intensive modeling. In contrast, generating 3D scenes from a single top-down image offers a lightweight and accessible alternative, making it possible to bootstrap comprehensive environments from minimal input.
@@
Despite its appeal, generating a coherent and complex 3D scene from a single image presents several fundamental challenges. First, the synthesized scene must exhibit consistent geometry across novel views, which is difficult for pure volumetric rendering methods such as Neural Radiance Fields (NeRF) (Mildenhall et al., 2021) or 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023). While these techniques excel at photorealistic appearance modeling, they often suffer from geometry artifacts, especially in occluded or sparsely visible regions, leading to multiview inconsistencies and structural implausibility (Li et al., 2022; Chung et al., 2023; Zhang et al., 2024a; Yu et al., 2024b). Second, the global layout of the generated scene must remain faithful to the input image. This is particularly challenging when considering the entire scene as a single asset and utilizing image-to-3D asset generators (Xiang et al., 2024; Team, 2025; Li et al., 2025), which often fail to preserve the spatial relationships between elements, resulting in distorted or semantically misaligned arrangements. Third, the local fidelity of individual objects should align closely with the visual evidence in the input. Due to the resolution constraints of 3D representations and the domain shift from object-level training to scene-level inference, previous image-to-3D generators are prone to producing low-quality meshes and misaligning textures.
@@
To address these challenges, we propose SceneFuse-3D, a training-free framework for generating complex 3D scenes from a single top-down image, by enhancing the capability of image-to-3D object generators. Our method combines two core components: region-based generation and spatial-aware 3D inpainting. Each targets specific challenges in existing pipelines. We divide the scene into overlapping regions and synthesize each independently. This modular approach enables spatial upscaling and improves local alignment by grounding generation on localized image crops. To maintain global coherence and object continuity across regions, we estimate a coarse 3D structure from monocular depth and landmark detection, forming a spatial prior. A masked rectified flow mechanism then completes missing parts while preserving known content, enhancing structural consistency and object-level fidelity throughout the scene.
@@
Through extensive experiments, we demonstrate that SceneFuse-3D generates realistic, diverse, and geometrically consistent 3D scenes from a single top-down image. Our method significantly outperforms strong baselines, including Trellis (Xiang et al., 2024), Hunyuan3D-2 (Team, 2025), TripoSG (Li et al., 2025), and LGM (Tang et al., 2024b), across both human preference and GPT-based evaluations. Quantitative results show notable gains in geometry quality, layout coherence, and texture fidelity, while qualitative comparisons highlight SceneFuse-3D’s ability to preserve spatial structure and fine-grained detail. These results underscore the effectiveness of our modular, training-free approach to 3D scene synthesis.
@@
  • We propose SceneFuse-3D, a training-free framework for generating structured 3D scenes from a single top-down image, leveraging pretrained object-centric generators for zero-shot scene asset synthesis.
@@
  • We conduct comprehensive evaluations on diverse scenes and show that SceneFuse-3D outperforms state-of-the-art baselines in geometry quality, layout coherence, and texture realism under both human and GPT-4o-based assessments.
@@ 2 Related Work
Progress in 2D generative models (Rombach et al., 2022; Ho et al., 2020; Chang et al., 2024; Ramesh et al., 2021) has enabled pipelines that outpaint views and then reconstruct 3D via depth fusion or volumetric representations such as NeRF (Mildenhall et al., 2021) and 3DGS (Kerbl et al., 2023). Early work focused on indoor scenes (Wiles et al., 2020; Koh et al., 2021; Höllein et al., 2023; Koh et al., 2023), with later methods tackling natural scenes (Li et al., 2022; Cai et al., 2023; Fridman et al., 2023; Chung et al., 2023) and improving reconstruction through stronger depth pipelines (Yu et al., 2024b; Zhang et al., 2024c; a; Yang et al., 2024; Shriram et al., 2024; Engstler et al., 2024; Yu et al., 2024a). Despite compelling renders, these approaches often exhibit geometric inconsistency due to hallucinations, especially in occluded regions. Panoramic (Stan et al., 2023; Li et al., 2024b; Wu et al., 2023; Schult et al., 2024; Liang et al., 2024) and multiview generation (Liu et al., 2024; Tang et al., 2023c; Gao et al., 2024) improve coverage but typically restrict camera motion and still struggle with accurate geometry. Concurrently, Syncity (Engstler et al., 2025) assembles block-wise 3D generations yet produces compact layouts and does not take full scene images as input. In contrast, we directly generate geometry-consistent scene assets with arbitrary layout from a single top-down image.
@@
Beyond 2D-to-3D pipelines, recent work directly generates scenes in native 3D. One line uses LLMs (Feng et al., 2023; Zhou et al., 2024; Çelen et al., 2024; Hu et al., 2024; Sun et al., 2023) or diffusion models (Tang et al., 2024a; Lin & Mu, 2024; Zhai et al., 2024; Vilesov et al., 2023; Maillard et al., 2024) to predict scene layouts that are then populated with assets. These methods yield semantically plausible arrangements but are often limited to predefined categories, struggle with coherent background geometry, and can suffer inter-object collisions. A separate line models scenes directly in latent 3D spaces (Wu et al., 2024; Meng et al., 2024; Ren et al., 2023; Liu et al., 2023b; Lee et al., 2024; Chai et al., 2023), e.g., BlockFusion’s triplane blocks (Wu et al., 2024) and LT3SD/XCube’s TUDF/voxel representations (Meng et al., 2024; Ren et al., 2023). While architecturally strong, these approaches require large, domain-specific 3D datasets (e.g., indoor rooms or urban layouts), limiting generalization to unseen scene types. In contrast, our method is training-free and modular, generating diverse 3D scene assets directly from a single top-down image.
@@
A parallel line of work generates single 3D assets from one image. Early methods combine 2D diffusion with NeRF/3DGS optimization (Poole et al., 2022; Lin et al., 2022; Tang et al., 2023b; Liu et al., 2023a; Tang et al., 2023a; 2024b) but often trade speed for geometry fidelity. Newer approaches directly generate 3D latents via diffusion (Gupta & Gupta, 2023; Xiong et al., 2024; Li et al., 2024a; Vahdat et al., 2022; Zhang et al., 2024b), and rectified-flow models further improve quality/efficiency (Xiang et al., 2024; Team, 2025; Li et al., 2025). However, these methods are trained on large object-centric datasets (e.g., Objaverse-XL (Deitke et al., 2023)), so applying them to full scenes faces limited 3D resolution and domain shift between objects and scenes, leading to spatial inconsistencies and layout hallucinations. We build on pretrained rectified-flow generators (Xiang et al., 2024) and mitigate these issues with a region-based strategy plus spatial-aware 3D inpainting, yielding scene assets with high geometric fidelity and global layout coherence from a single top-down image.
@@ 3 Method
Figure 2: Overview of the SceneFuse-3D Pipeline. Given a single top-down image, we first estimate a coarse scene structure via monocular depth and landmark extraction to initialize the scene latent (Spatial Prior Initialization). The scene is divided into overlapping regions for localized synthesis and progressively fused into a coherent global latent (Region-based Generation & Fusion). Each region is completed using a two-stage masked rectified flow pipeline with a sparse structure generator $\mathcal{G}_{s}$ and a structured latent generator $\mathcal{G}_{L}$ (Spatial-aware 3D Completion). The final 3D scene is decoded from the completed structured latent.
@@
From the scene image, we first construct the scene latents with structured latent representations (Sec. 3.1) from spatial priors (Sec. 3.2). Then, we divide the scene-level latent into region-level latents for sequential processing. For each region, we extract the latents from the latest scene-level structured latents and take those as priors for spatial-aware 3D completion (Sec. 3.3) by using the base 3D generator, Trellis (Xiang et al., 2024). Later, we fuse the updated region latents to the scene latents for cross-region consistency (Sec. 3.4). After finishing all regions, we leverage pretrained object decoders on the complete scene structured latents to obtain a 3D scene asset. Details are explained below.
@@ 3.1 Structured Latent Representation
To leverage the pretrained knowledge of the base 3D generator, we construct the scene with a structured latent representation $z$ (Xiang et al., 2024):
@@
For the general image-to-3D asset generation, there are two pretrained rectified flow transformers: a sparse structure generator $\mathcal{G}_{s}$ and a structured latent generator $\mathcal{G}_{L}$. At inference time, $\mathcal{G}_{s}$ first generates the active voxel positions $\{p_{i}\}^{L}$ from the noisy grids $V_{T}$ with Gaussian noise. These positions are then used by $\mathcal{G}_{L}$ to generate the corresponding latent features $\{f_{i}\}^{L}$ from noisy features $F_{T}$. Both generators are conditioned on an input image condition $C_{I}$, encoded by DINOv2 (Oquab et al., 2023). Finally, the structured latent representation $z$ is decoded into a 3D object $O$ using object decoders, which include different sparse 3D VAE decoders for 3D Gaussians, Radiance Fields, and mesh generation. The overall process is described as follows:
@@ 3.2 Spatial Prior Initialization
Given a top-down image, the image-to-3D generator may produce outputs in arbitrary orientations. To resolve this ambiguity and provide a consistent structural prior, we initialize the scene using point clouds. Specifically, we employ a monocular depth estimator (Wang et al., 2025) to predict a depth image and infer camera parameters, from which we construct pixel-wise point clouds. Due to occlusions, these point clouds contain missing regions, which will later be filled by the image-to-3D generator.
However, occluded areas of complex objects (e.g., buildings) may have multiple plausible completions. To enforce cross-region consistency in such cases, we propose first generating landmark objects independently and then conditioning subsequent generations on their geometry. To achieve this, we use Florence2 (Xiao et al., 2024) to propose landmark bounding boxes and SAM2 (Ravi et al., 2024) to extract instance masks. Each detected landmark is then processed individually using the 3D generator to obtain instance-level meshes. These meshes are aligned with the raw point clouds using ICP (Rusinkiewicz & Levoy, 2001), replacing the original landmark regions with mesh-derived point clouds.
@@ 3.3 Region-based Generation
While region-based generation improves local fidelity, it introduces a new challenge: how to maintain global consistency, especially across overlapping regions. Furthermore, since the 3D generator may create assets from any orientation, we need to alleviate the misalignment between image conditions and region latents to enable further region fusion. To address these challenges, we draw inspiration from training-free inpainting methods in 2D diffusion models, such as RePaint (Lugmayr et al., 2022), and adapt a similar approach for 3D generation. We propose to use a masked rectified flow pipeline for 3D completion, which treats the partially completed global scene latent as a constraint and performs conditional generation over the current region by completing only the unknown parts.
@@
We initialize the latent variable $x_{T}\sim\mathcal{N}(0,I)$ with Gaussian noise, representing the unknown region at the final time step. For each timestep $t=T,T{-}1,\dots,1$, we perform $U$ resampling steps to improve stability and smoothness (Lugmayr et al., 2022). At each resampling iteration $u$, we first compute the flow field $v_{\theta}(x_{t},t)$ using the rectified flow model and apply an Euler update to obtain the intermediate latent:
@@
$\texttt{forward\_step}(x,t)=(1-t)\cdot x+\left[\sigma_{\min}+(1-\sigma_{\min})\cdot t\right]\cdot\epsilon,\quad\epsilon\sim\mathcal{N}(0,I),$ (10)